A wheel is accelerated from rest to 30 rad/s by torque of 200 Nm. The wheel turns

through 20 radians during this time.

Calculate the:

i. Angular acceleration


1
Expert's answer
2021-12-27T08:00:46-0500

"\phi =\frac{\omega^2-\omega_0^2}{2\epsilon}\to\epsilon=\frac{\omega^2-\omega_0^2}{2\phi}=\frac{30^2-0}{2\cdot20}=22.5\ (rad/s^2)"